Which

Show the full path of commands

Edition 2.20, for Which Version 2.20 16 January 2008

Carlo Wood <carlo@gnu.org>

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000 - 2007, by Carlo Wood, Run on IRC <carlo@alinoe.com> RSA-1024 0x624ACAD5 1997-01-26 Sign & Encrypt Fingerprint16 = 32 EC A7 B6 AC DB 65 A6 F6 F6 55 DD 1C DC FF 61

1 The which Program

Which takes one or more arguments. For each of its arguments it prints to stdout the full path of the executables that would have been executed when this argument had been entered at the shell prompt. It does this by searching for an executable or script in the directories listed in the environment variable PATH using the same algorithm as bash(1).

2 which: show full path of commands Chapter 2: Invoking which 3

2 Invoking which

The synopsis to invoke which is

     which [options] [--] programname [...]

4 which: show full path of commands

3 Option Summary

‘--all’

‘-a’ Print all matching executables in PATH, not just the first.

‘--read-alias’

‘-i’ Read aliases from stdin, reporting matching ones on stdout. This is useful in combination with using an alias for which itself. For example

alias which='alias | which -i'.
‘--skip-alias’

Ignore option ‘--read-alias’, if any. This is useful to explicity search for normal binaries, while using the ‘--read-alias’ option in an alias or function for which.

‘--read-functions’

Read shell function definitions from stdin, reporting matching ones on stdout. This is useful in combination with using a shell function for which itself. For example:

which() { declare -f | which --read-functions $@ }
export -f which
‘--skip-functions’

Ignore option ‘--read-functions’, if any. This is useful to explicity search for normal binaries, while using the ‘--read-functions’ option in an alias or function for which.

‘--skip-dot’

Skip directories in PATH that start with a dot.

‘--skip-tilde’

Skip directories in PATH that start with a tilde and executables which reside in the HOME directory.

‘--show-dot’

If a directory in PATH starts with a dot and a matching executable was found for that path, then print "./programname" rather than the full path.

‘--show-tilde’

Output a tilde when a directory matches the HOME directory. This option is ignored when which is invoked as root.

‘--tty-only’

Stop processing options on the right if not on tty.

‘--version, -v, -V’

Print version information on standard output then exit successfully.

‘--help’ Print usage information on standard output then exit successfully.

6 which: show full path of commands Chapter 4: Return Value 7

4 Return Value

Which returns the number of failed arguments, or -1 when no ‘programname’ was given.

8 which: show full path of commands

5 Example

The recommended way to use this utility is by adding an alias (C shell) or shell function (Bourne shell) for which like the following:

[ba]sh:

which ()
{
  (alias; declare -f) | /usr/bin/which --tty-only --read-alias --read-functions --show-tilde --show-dot $@
}
export -f which

[t]csh:

alias which 'alias | /usr/bin/which --tty-only --read-alias --show-dot --show-tilde'

This will print the readable ~/ and ./ when starting which from your prompt, while still printing the full path when used from a script:

> which q2
~/bin/q2
> echo `which q2`
/home/carlo/bin/q2

10 which: show full path of commands

6 Bugs

The HOME directory is determined by looking for the HOME environment variable, which aborts when this variable doesn’t exist. Which will consider two equivalent directories to be different when one of them contains a path with a symbolic link.

12 which: show full path of commands Chapter 7: See Also 13

7 See Also^{bash(1)}

14 which: show full path of commands Index 15

Index

-

‘--all’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--help’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--read-alias’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--read-functions’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--show-dot’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--show-tilde’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--skip-alias’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--skip-dot’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--skip-functions’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--skip-tilde’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--tty-only’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘--version’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘-a’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘-i’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘-v’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 ‘-V’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

A

aliases, handling of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

B

Bugs, known . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

C

Command line Options . . . . . . . . . . . . . . . . . . . . . . . . . 5

D

Description of which . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

E

Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

I

Invoking which . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

O

Options, command line . . . . . . . . . . . . . . . . . . . . . . . . . 5 Overview of command line options . . . . . . . . . . . . . . 5

R

Return value of which . . . . . . . . . . . . . . . . . . . . . . . . . . 7

S

Synopsis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

16 which: show full path of commands

Table of Contents

1 The which Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Invoking which . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Option Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Return Value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 5 Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6 Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 7 See Also . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

ii which: show full path of commands